OpenStudy (anonymous):

Find the Maclaurin Series as far as the fourth term for the following equation: cos^2 2x

OpenStudy (anonymous):

f(x) = f(0) + f'(0)x + f''(0)/2! x^2 +..... = cos^2 2)0) + d(cos^2 2x) dx *x +.. sorry gotta go

OpenStudy (anonymous):

d(cos^2 2x) dx *x = -4 sin^2 2x*x

OpenStudy (anonymous):

= -4x sin^2 2x - sorry i have to go

OpenStudy (anonymous):

okay... im just trying to understand what you've done. I think i've got my head around it

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=Taylor+Cos^2+%282x%29 Better check it.....

OpenStudy (anonymous):

but thats taylor series

OpenStudy (anonymous):

Taylor about 0 = MacLaurin

OpenStudy (anonymous):

so thats the correct answer

OpenStudy (anonymous):

Wolfram says it is, do it manually if u don't believe it...